Please help me with these two questions!!!!! I’m really lost!

1. The neurotransmitter measured by an electrode is dopamine. Each dopamine molecule that diffuses to the electrode releases two electrons. The charge transferred to the electrode by one burst is 0.13 pC (picocolombs). One coulomb of charge corresponds to 6.24 x 10^18 electrons. How many molecules are released in one burst?

2. A 120-pound woman working in an office expends about 2.2 x 103 kcal/day, whereas the same woman climbing a mountain needs 3.4 x 103 kcal/day.    A) how many joules per day does the woman expend in each activity?

B) how many watts does the woman expend in each activity? Which consumers more power, the office worker or a 100-W light bulb?
Answer #1

1. 1pC = 10^-12 C , then 0.13pC = 1.3×10^-13 C

As 1 C = 6.24×10^18 electrons

Then , 1.3×10^-13 C would produced by = (6.24×10^18) × (1.3×10^-13) = 8.112×10^5 electrons.

And as 2 electrons are released by = 1 molecule

Then , 8.112×10^5 electrons would released by = (8.112×10^5)÷2 = 4.06×10^5 molecules.

Thus 4.06×10^5 molecules must released in one burst.